# An additive two-level parallel variant of the DMRG algorithm with coarse-space correction

### v1

@@ A.2. Computational Scaling of the A2DMRG Algorithm 2
Let us now consider the computational cost of the additive variant of the two-site DMRG algorithm. Mimicking the same arguments as those used to study the one-site DMRG case, we see that the computational cost of the first two steps of Algorithm (2) are given by
[…]
The third step in the additive two-site DMRG algorithm requires the computation of the lowest eigenvalue of the symmetric eigenvalue problem (18). As in the one-site DMRG algorithm case, this can either be done using an explicit construction of the underlying matrices $\widehat{\mathbb{S}}\in\mathbb{R}^{d\times d}$ and $\widehat{\mathbb{A}}$ or by relying on a Krylov solver. To evaluate the cost of the latter approach, it suffices to check that a factorization of the form (36)-(37) can also be obtained for linear combinations of tensor train decompositions obtained from the second step of the additive two-site DMRG algorithm 2. We claim that this is indeed the case. Indeed, consider a linear combination of tensor of the form
$\sum_{j=1}^{d}{\mathbb{y}}_{j}\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)}),$
where each ${\mathbb{y}}_{j}\in\mathbb{R}$ and each tensor train decomposition $\widetilde{\mathbb{Y}}_{j}^{(n+1)}$ is obtained as the solution to the $j^{\rm th}$ two-site DMRG micro-step in accordance with step two of the A2DMRG Algorithm 2. By construction, each TT decomposition $\widetilde{\mathbb{Y}}_{j}^{(n+1)}$ takes the form
$\widetilde{\mathbb{Y}}_{j}^{(n+1)}=\big{(}\mathbb{U}_{1}^{(n)},\ldots\mathbb{U}_{j-1}^{(n)},\delta\mathbb{U}_{j}^{(n+1)},\delta\mathbb{U}_{j+1}^{(n+1)},\mathbb{V}_{j+2}^{(n)},\ldots\mathbb{V}_{d}^{(n)}\big{)}$
where the components $\delta\mathbb{U}_{j}^{(n+1)}\in\mathbb{R}^{r_{j-1}\times n_{j}\times\widetilde{r}_{j}}$ and $\delta\mathbb{U}_{j+1}^{(n+1)}\in\mathbb{R}^{\widetilde{r}_{j}\times n_{j}\times r_{j+1}}$ have been obtained through the $j^{\rm th}$ two-site DMRG micro-iteration applied to the initial $j$-orthogonal tensor train iterate $\mathbb{Y}_{j}^{(n)}$ followed by a truncated SVD step. It now follows from a direct calculation that for all $x_{j}\in\{1,\ldots,n_{j}\},~{}j=1,\ldots,d$ we can write
(41) $\Big{(}\sum_{j=1}^{d}\mathbb{t}_{j}\tau(\widetilde{Y}_{j}^{(n+1)})\Big{)}(x_{1},\ldots,x_{d})$ $=\mathbb{W}_{1}(x_{1},x_{2})\mathbb{W}_{2}(x_{2},x_{3})\ldots\mathbb{W}_{d-1}(x_{d-1},x_{d}),$
@@
(42) $\begin{split}\mathbb{W}_{1}(x_{1},x_{2})&=\begin{bmatrix}\mathbb{t}_{1}\delta\mathbb{U}^{(n+1)}_{1}(x_{1})\delta\mathbb{U}^{(n+1)}_{2}(x_{2})&\mathbb{U}_{1}^{(n)}(x_{1})\end{bmatrix}\in\mathbb{R}^{r_{0}\times(r_{2}+r_{1})},\\[10.00002pt] \mathbb{W}_{j}(x_{j},x_{j+1})&=\begin{bmatrix}\mathbb{V}^{(n+1)}_{j+1}(x_{j+1})&0\\[3.99994pt] \mathbb{t}_{j}\delta\mathbb{U}^{(n+1)}_{j}(x_{j})\delta\mathbb{U}^{(n+1)}_{j+1}(x_{j+1})&\mathbb{U}^{(n+1)}_{j}(x_{j})\\ \end{bmatrix}\in\mathbb{R}^{(r_{j}+r_{j-1})\times(r_{j+1}+r_{j})},\\ \noindent&\hskip 213.39566pt\forall j\in\{2,\ldots,d-2\},\\[10.00002pt] \mathbb{W}_{d-1}(x_{d-1},x_{d})&=\begin{bmatrix}\mathbb{V}^{(n+1)}_{d}(x_{d})\\[3.99994pt] \mathbb{t}_{d}\delta\mathbb{U}_{d-1}^{(n+1)}(x_{d-1})\delta\mathbb{U}_{d}^{(n+1)}(x_{d})\end{bmatrix}\in\mathbb{R}^{(r_{d-1}+r_{d-2})\times r_{d}}.\end{split}$
@@
Although Equation (41) does not correspond to a tensor train decomposition of the tensor $\sum_{j=1}^{d}\mathbb{t}_{j}\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)})$, the structure of this factorization closely resembles a tensor train decomposition, and we can therefore use the usual TT-type tensor contractions to efficiently compute inner products involving tensors of the form $\sum_{j=1}^{d}\mathbb{t}_{j}\tau(\widetilde{Y}_{j}^{(n+1)})$. In particular, the computational cost of calculating an inner-product of the form (c.f., Equation (35))
$\left\langle\sum_{j=1}^{d}{\mathbb{t}}_{j}\widetilde{Y}_{j}^{(n+1)},\mathbb{A}\sum_{k=1}^{d}\mathbb{w}_{k}\mathbb{y}_{\ell}\widetilde{Y}_{k}^{(n+1)}\right\rangle,$
@@
for arbitrary vectors $\mathbb{t},\mathbb{w}\in\mathbb{R}^{d}$ scales as $\mathcal{O}(d^{2}n^{3}r^{3}R+dn^{2}\widetilde{r}r^{2})$, where we have introduced $\widetilde{r}=\max\{\widetilde{r}_{j}\}_{j=1}^{d-1}$. Assuming therefore, that $K^{\prime}\in\mathbb{N}$ Krylov iterations are required to solve the eigenvalue problem (18), we arrive at the following total computational cost of this third step of the additive two-site DMRG algorithm 2:
[…]
$=$ $\mathcal{O}\big{(}d^{2}n^{3}r^{3}RK^{\prime}+dn^{2}\widetilde{r}r^{2}\big{)}+\mathcal{O}(d^{3}n\widetilde{r}^{3}+d^{3}).$
It remains to consider the computational cost of the fourth and final step in the additive two-site DMRG algorithm, namely the TT rounding step. In contrast to the one-site DMRG case, an additional complication now arises, namely, that the factorization (41)-(42) of a linear combination of tensor train decompositions does not itself represent a tensor train decomposition of ranks bounded by $2\mathbb{r}$. Consequently, we can no longer appeal directly to the TT rounding algorithm of Oseledets [33]. Instead, we propose to rephrase the TT rounding step as a best-approximation problem on the manifold $\mathcal{T}_{\widetilde{r}}\subset\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ with ranks $(\widetilde{r}_{1},\ldots,\widetilde{r}_{d-1})$. In other words, given the solution $\sum_{j=1}^{d}{\mathbb{c}}_{j}^{*}\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)})$ to the eigenvalue problem (18), we consider the minimization problem
(43) $\underset{\mathbb{X}\in\mathcal{T}_{\widetilde{\mathbb{r}}}}{\text{argmin}}\frac{1}{2}\|\sum_{j=1}^{d}{\mathbb{c}}_{j}^{*}\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)})-\mathbb{X}\|^{2}=\underset{\mathbb{X}\in\mathcal{T}_{\widetilde{\mathbb{r}}}}{\text{argmin}}\frac{1}{2}\langle\mathbb{X},\mathbb{X}\rangle-\Big{\langle}\sum_{j=1}^{d}{\mathbb{c}}_{j}^{*}\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)}),\mathbb{X}\Big{\rangle}.$
Equation (43) can, for instance, be solved by making use of the one-site A2DMRG Algorithm 2 together with the factorization (41)-(42). Recalling the computational cost of each iteration of the one-site A2DMRG algorithm Algorithm 2 given by Equation (38), and denoting by $M\in\mathbb{N}$ the total number of global iterations required for convergence up to a tolerance, this step scales as
[…]
We thus conclude that each global iteration of the additive algorithm 2 has computational cost
[…]
$=\mathcal{O}\big{(}dnr^{3}R(Kn+d)\big{)}+\mathcal{O}\big{(}d^{2}n^{3}r^{3}RK^{\prime}+dn^{2}\widetilde{r}r^{2}\big{)}+\mathcal{O}(d^{3}n\widetilde{r}^{3})$
@@
(45) $+\mathcal{O}\big{(}dn\widetilde{r}^{3}(K_{\rm 1Site}+d)M\big{)}+\mathcal{O}\big{(}d^{2}n\widetilde{r}^{3}K_{\rm 1Site}^{\prime}\big{)}.$
@@
${\rm Cost}_{\rm 2site,compute}^{\rm add,pp}$ $=\mathcal{O}\big{(}nr^{3}R(Kn+d)\big{)}+\mathcal{O}\big{(}dn^{3}r^{3}RK^{\prime}+n^{2}\widetilde{r}r^{2}\big{)}+\mathcal{O}(d^{2}n\widetilde{r}^{3})$
$+\mathcal{O}\big{(}n\widetilde{r}^{3}(K_{\rm 1Site}+d)M\big{)}+\mathcal{O}\big{(}dn\widetilde{r}^{3}K_{\rm 1Site}^{\prime}\big{)}.$

### v2

@@ A.2. Computational Scaling of the A2DMRG Algorithm 2
Let us now consider the computational cost of the two-site A2DMRG algorithm. Mimicking the same arguments as those used to study the one-site DMRG case, we see that the computational cost of the first two steps of Algorithm (2) are given by
${\rm Cost}^{\rm 2site}_{\rm Step\;1}+{\rm Cost}^{\rm 2site}_{\rm Step\;2}=\mathcal{O}\big(dnr^{3}\big)+\mathcal{O}\big(dnr^{3}R(nK+d)\big).$
The third step in the two-site A2DMRG algorithm requires the computation of the lowest eigenvalue of the symmetric eigenvalue problem (18). As in the one-site DMRG algorithm case, this can either be done using an explicit construction of the underlying matrices $\widehat{\mathbb{S}}\in\mathbb{R}^{d\times d}$ and $\widehat{\mathbb{A}}$ or by relying on a Krylov solver. To evaluate the cost of the latter approach, it suffices to check that a factorization of the form (36)-(37) can also be obtained for linear combinations of tensor train decompositions obtained from the second step of the two-site A2DMRG Algorithm 2. We claim that this is indeed the case. Indeed, consider a linear combination of tensor of the form
$\sum_{j=1}^{d}{\mathbb{y}}_{j}\tau(\widetilde{\mathbb{U}}^{(n+1),{j-1}}),$
where each ${\mathbb{y}}_{j}\in\mathbb{R}$, and each tensor train decomposition $\widetilde{\mathbb{U}}^{(n+1),j},~j\in\{1,\ldots,d-1\}$ is obtained as the solution to the $j^{\rm th}$ two-site DMRG micro-step in accordance with step two of the A2DMRG Algorithm 2. By construction, each TT decomposition $\widetilde{\mathbb{U}}^{(n+1),j},~j\in\{1,\ldots,d-1\}$ takes the form
$\widetilde{\mathbb{U}}^{(n+1),j}=\big(\mathbb{U}_{1}^{(n)},\ldots\mathbb{U}_{j-1}^{(n)},\delta\mathbb{U}_{j}^{(n+1)},\delta\mathbb{U}_{j+1}^{(n+1)},\mathbb{V}_{j+2}^{(n)},\ldots\mathbb{V}_{d}^{(n)}\big)$
where the components $\delta\mathbb{U}_{j}^{(n+1)}\in\mathbb{R}^{r_{j-1}\times n_{j}\times\widetilde{r}_{j}}$ and $\delta\mathbb{U}_{j+1}^{(n+1)}\in\mathbb{R}^{\widetilde{r}_{j}\times n_{j}\times r_{j+1}}$ have been obtained through the $j^{\rm th}$ two-site DMRG micro-iteration applied to the initial $j$-orthogonal tensor train iterate $\mathbb{U}^{(n),j}$ followed by a truncated SVD step. Since $\tau({\mathbb{U}}^{(n+1),0})=\tau({\mathbb{U}}^{(n),j})$ for all $j\in\{1,\ldots,d-1\}$, we can once again take the right-orthogonal tensor train decomposition of $\tau({\mathbb{U}}^{(n+1),0})$, i.e.,
$\tau({\mathbb{U}}^{(n+1),0})=\tau\big({\delta W_{1}}^{(n)},\mathbb{V}_{2}^{(n)}\ldots\mathbb{V}_{d}^{(n)}\big).$
It now follows from a direct calculation that for all $x_{\ell}\in\{1,\ldots,n_{\ell}\},~\ell=1,\ldots,d$ we can write
(41) $\Big(\sum_{j=1}^{d}\mathbb{y}_{j}\tau(\widetilde{\mathbb{U}}^{(n+1),{j-1}})\Big)(x_{1},\ldots,x_{d})$ $=\mathbb{W}_{1}(x_{1},x_{2})\mathbb{W}_{2}(x_{2},x_{3})\ldots\mathbb{W}_{d-1}(x_{d-1},x_{d}),$
@@
(42) $\begin{split}\mathbb{W}_{1}(x_{1},x_{2})&=\begin{bmatrix}\mathbb{y}_{2}\delta\mathbb{U}^{(n+1)}_{1}(x_{1})\delta\mathbb{U}^{(n+1)}_{2}(x_{2})+\mathbb{y}_{1}\delta\mathbb{W}^{(n)}_{1}(x_{1})\mathbb{V}^{(n)}_{2}(x_{2})&\mathbb{U}_{1}^{(n)}(x_{1})\end{bmatrix}\in\mathbb{R}^{r_{0}\times(r_{2}+r_{1})},\\[10.00002pt] \mathbb{W}_{\ell}(x_{\ell},x_{\ell+1})&=\begin{bmatrix}\mathbb{V}^{(n+1)}_{\ell+1}(x_{\ell+1})&0\\[3.99994pt] \mathbb{y}_{\ell+1}\delta\mathbb{U}^{(n+1)}_{\ell}(x_{\ell})\delta\mathbb{U}^{(n+1)}_{\ell+1}(x_{\ell+1})&\mathbb{U}^{(n+1)}_{\ell}(x_{\ell})\\ \end{bmatrix}\in\mathbb{R}^{(r_{\ell}+r_{\ell-1})\times(r_{\ell+1}+r_{\ell})},\\ \noindent&\hskip 213.39566pt\forall\ell\in\{2,\ldots,d-2\},\\[10.00002pt] \mathbb{W}_{d-1}(x_{d-1},x_{d})&=\begin{bmatrix}\mathbb{V}^{(n+1)}_{d}(x_{d})\\[3.99994pt] \mathbb{y}_{d}\delta\mathbb{U}_{d-1}^{(n+1)}(x_{d-1})\delta\mathbb{U}_{d}^{(n+1)}(x_{d})\end{bmatrix}\in\mathbb{R}^{(r_{d-1}+r_{d-2})\times r_{d}}.\end{split}$
@@
Although Equation (41) does not correspond to a tensor train decomposition of the tensor $\sum_{j=1}^{d}\mathbb{y}_{j}\tau(\widetilde{\mathbb{U}}^{(n+1),{j-1}})$, the structure of this factorization closely resembles a tensor train decomposition, and we can therefore use the usual TT-type tensor contractions to efficiently compute inner products involving tensors of the form $\sum_{j=1}^{d}\mathbb{y}_{j}\tau(\widetilde{\mathbb{U}}^{(n+1),{j-1}})$. In particular, the computational cost of calculating an inner-product of the form (c.f., Equation (35))
$\left\langle\sum_{j=1}^{d}{\mathbb{t}}^{(i)}_{j}\widetilde{\mathbb{U}}^{(n+1),{j-1}},\mathbb{A}\sum_{k=1}^{d}\mathbb{w}_{k}\widetilde{\mathbb{U}}^{(n+1),{k-1}}\right\rangle,$
@@
for arbitrary vectors $\mathbb{t}^{(i)},\mathbb{w}\in\mathbb{R}^{d}$ scales as $\mathcal{O}(d^{2}n^{3}r^{3}R+dn^{2}\widetilde{r}r^{2})$, where we have introduced $\widetilde{r}=\max\{\widetilde{r}_{j}\}_{j=1}^{d-1}$. Assuming therefore, that $K^{\prime}\in\mathbb{N}$ Krylov iterations are required to solve the eigenvalue problem (18), we arrive at the following total computational cost of this third step of the two-site A2DMRG Algorithm 2:
@@
${\rm Cost}_{\rm Step\;3,Krylov}^{\rm 2site}=$ $\mathcal{O}\big(d^{2}n^{3}r^{3}RK^{\prime}+dn^{2}\widetilde{r}r^{2}\big)+{\rm Cost}_{\mathbb{S},\rm SVD}$
$=$ $\mathcal{O}\big(d^{2}n^{3}r^{3}RK^{\prime}+dn^{2}\widetilde{r}r^{2}K^{\prime}\big)+\mathcal{O}(d^{3}n\widetilde{r}^{3}+d^{3}).$
It remains to consider the computational cost of the fourth and final step in the two-site A2DMRG algorithm, namely the TT rounding step. In contrast to the one-site DMRG case, an additional complication now arises, namely, that the factorization (41)-(42) of a linear combination of tensor train decompositions does not itself represent a tensor train decomposition of ranks bounded by $2\mathbb{r}$. Consequently, we can no longer appeal directly to the TT rounding algorithm of Oseledets [35]. Instead, we propose to rephrase the TT rounding step as a best-approximation problem on the manifold $\mathcal{T}_{\widetilde{r}}\subset\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ with ranks $(\widetilde{r}_{1},\ldots,\widetilde{r}_{d-1})$. In other words, given the solution $\sum_{j=0}^{d}{\mathbb{c}}_{j}^{*}\tau(\widetilde{\mathbb{U}}^{(n+1),{j}})$ to the eigenvalue problem (18), we consider the minimization problem
(43) $\underset{\mathbb{X}\in\mathcal{T}_{\widetilde{\mathbb{r}}}}{\text{argmin}}\frac{1}{2}\|\sum_{j=0}^{d}{\mathbb{c}}_{j}^{*}\tau(\widetilde{\mathbb{U}}^{(n+1),j})-\mathbb{X}\|^{2}=\underset{\mathbb{X}\in\mathcal{T}_{\widetilde{\mathbb{r}}}}{\text{argmin}}\frac{1}{2}\langle\mathbb{X},\mathbb{X}\rangle-\Big\langle\sum_{j=0}^{d}{\mathbb{c}}_{j}^{*}\tau(\widetilde{\mathbb{U}}^{(n+1),j}),\mathbb{X}\Big\rangle.$
Equation (43) can, for instance, be solved using the one-site A2DMRG Algorithm 2 together with the factorization (41)-(42). Recalling the computational cost of each iteration of the one-site A2DMRG algorithm Algorithm 2 given by Equation (38), and denoting by $M^{\prime}\in\mathbb{N}$ the total number of global iterations required for convergence up to a tolerance, this step scales as
(44) ${\rm Cost}_{\rm Step\;4}^{\rm 2site}=\mathcal{O}\big(dn\widetilde{r}^{3}(K_{\rm 1Site}+d)M^{\prime}\big)+\mathcal{O}\big(d^{2}n\widetilde{r}^{3}K_{\rm 1Site}^{\prime}M^{\prime}\big)+\mathcal{O}(M^{\prime}d^{3}),$
where $K_{\rm 1Site},K_{\rm 1Site}^{\prime}$ denote the number of Krylov solver iterations required for the one-site DMRG micro-iterations and the solution to the second-level minimization problem (18) respectively. We emphasize here that the application of the one-site A2DMRG algorithm to the best-approximation problem (43) requires the solution of a linear system of equations, both for the micro-iterations and the second-level minimization. Consequently, the computational cost ${\rm Cost}_{\rm Step\;4}^{\rm 2site}$ given by Equation (44) differs slightly from the computational cost of one-site A2DMRG calculated earlier (see Equation (38)).
We thus conclude that each global iteration of the A2DMRG Algorithm 2 has computational cost
${\rm Cost}_{\rm 2site,compute}^{\rm A2DRMG}$ $={\rm Cost}^{\rm 2site}_{\rm Step\;1}+{\rm Cost}^{\rm 2site}_{\rm Step\;2}+{\rm Cost}^{\rm 2site}_{\rm Step\;3,Krylov}+{\rm Cost}^{\rm 2site}_{\rm Step\;4}$
$=\mathcal{O}\big(dnr^{3}R(Kn+d)\big)+\mathcal{O}\big(d^{2}n^{3}r^{3}RK^{\prime}+dn^{2}\widetilde{r}r^{2}K^{\prime}\big)+\mathcal{O}(d^{3}n\widetilde{r}^{3})$
(45) $+\mathcal{O}\big(dn\widetilde{r}^{3}(K_{\rm 1Site}+d)M^{\prime}\big)+\mathcal{O}\big(d^{2}n\widetilde{r}^{3}K_{\rm 1Site}^{\prime}M^{\prime}\big).$
@@
${\rm Cost}_{\rm 2site,compute}^{\rm A2DRMG,pp}$ $=\mathcal{O}\big(nr^{3}R(Kn+d)\big)+\mathcal{O}\big(dn^{3}r^{3}RK^{\prime}+n^{2}\widetilde{r}r^{2}K^{\prime}\big)+\mathcal{O}(d^{2}n\widetilde{r}^{3})$
$+\mathcal{O}\big(n\widetilde{r}^{3}(K_{\rm 1Site}+d)M^{\prime}\big)+\mathcal{O}\big(dn\widetilde{r}^{3}K_{\rm 1Site}^{\prime}M^{\prime}\big).$
Let us remark that throughout our testing, $K_{\rm 1site},K_{\rm 1site}^{\prime}\approx\mathcal{O}(1)$ so that the last two terms are dominated in cost by the first two terms.